Which measure of central tendency do you think best represents the central tendency of the data? 1, 2, 2, 2, 2, 3, 3, 3, 3, 4, 4, 4, 9, 1850
Answer #1

The given data set is: 1, 2, 2, 2, 2, 3, 3, 3, 3, 4, 4, 4, 9, 1850

Here, one of the data value is a large outlier. If there is any outlier, that can influence the mean heavily. So, it is not suitable to use mean as a measure of central tendency. Therefore, median can be considered as the best suited measure of central tendency.

Number of data values = 14

Median = (14+1)/2 = 7.5 th value

That is, average of 7th and 8th value = (3+3)/2 = 3